# QCD evolution of entanglement entropy

### v1

@@ I Introduction
Quantum entanglement, famously described by Albert Einstein as a ”spooky action at a distance,” is fundamental to our understanding of the microscopic world. The nontrivial and non-local correlations among the constituents of a quantum system can provide deep insights into the underlying properties of such systems. In recent years, there has been a growing interest in quantum entanglement within high-energy physics, particularly in understanding the structure of hadrons Kharzeev and Levin (2017), thermalization Alba and Calabrese (2017) (see Berges et al. (2021) for a review), spin and helicity coupling of partons Bhattacharya et al. (2024), violation of Bell inequalities in hadronic final states Gong et al. (2022); Aoude et al. (2022); Severi et al. (2022); White and White (2024); CMS (2024); Aad et al. (2023); Banuls et al. (2023), and confinement Klebanov et al. (2008); Grieninger et al. (2023a).
@@
The state produced from the proton after the DIS measurement is therefore a mixed state. The corresponding density matrix is obtained by tracing the density matrix of the pure proton state over the unobserved phases of the Fock states, and is characterized by non-zero von Neumann entropy. This entropy can be viewed as the entropy of entanglement between the occupation numbers and the corresponding phases in the Fock state decomposition of the proton wave function. Once this entropy is produced, it will be observed in the final state through the entropy of the multi-hadron state created in DIS. In general, the entanglement entropy may differ from the entropy of the final hadronic state. However there are several examples where the produced entanglement entropy appears equal to the entropy of the final state. One can thus hypothesize that the entanglement entropy is equal to the entropy of the hadron state, and moreover that the hadron multiplicity distribution mirrors the parton occupation number distribution inside the proton. This hypothesis can be viewed as a stronger form of the “local parton-hadron duality”. In previous studies, it has been found that this bold assumption holds surprisingly well Tu et al. (2020); Hentschinski and Kutak (2022); Hentschinski et al. (2022, 2023). Interestingly the model predicts that the entropy linearly grows with rapidity what is in accord with consideration based of near conformal invariance of QCD in the high energy limit Lipatov (1986, 1990, 1994); Braun et al. (2003) and the symmetries of the considered dipole model Gürsoy et al. (2023); Caputa and Kutak (2024). For other works on entanglement entropy at the high energy regime, see Kutak (2011); Peschanski (2013); Armesto et al. (2019); Neill and Waalewijn (2019); Armesto et al. (2019); Kovner et al. (2019); Chachamis et al. (2024); Liu et al. (2022a, 2023a, b, 2023b); Stoffers and Zahed (2013); Asadi and Vaidya (2023); Kou et al. (2022); Kutak (2023); Dumitru et al. (2023); Ehlers (2022, 2023); Florio et al. (2024); Grieninger et al. (2024, 2023b); Ikeda et al. (2023); Brandenburg et al. (2024); Berges et al. (2018a, b); Dumitru and Kolbusz (2023, 2022); Ramos and Machado (2022); Moriggi et al. (2024); Ramos and Machado (2020).
@@ II Entanglement entropy in inclusive DIS
In the following we explore both versions of entanglement entropy. For the universal entanglement entropy
@@
Here $g(x,\mu)$ and $q_{f}(x,\mu)$ denote the PDF of the gluon and the quark of flavor $f$, respectively, which depend on the factorization scale which for the DIS reaction is usually identified with the photon virtuality $Q$. The $\mu$ dependence of PDFs is obtained from the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution equations Gribov and Lipatov (1972); Altarelli and Parisi (1977); Dokshitzer (1977).
@@ II.1 Definition of a local dipole probability distribution
The probability to have $n$ dipoles in the interval $[0,y]$ is given by all possible combinations to have $m$ dipoles in the range $[0,y_{0}]$ and $n-m$ in the range $[y_{0},y]$. With
@@ II.2 Definition of local entropy
with a $C_{0}\leq 1$ a constant factor which accounts for subleading effects that modify the no emission probability. With this parameter we will have in particular $\tilde{p}_{0}(y,y)=C_{0}\neq 1$, i.e. the probability for emission, $1-\tilde{p}_{0}$, is non-zero if the rapidity range $y-y_{0}$ shrinks to zero. It therefore introduces a certain non-locality in the parton-hadron duality, which relates the generated dipoles to final state hadrons. While the overall number of dipoles is fixed by the PDFs, a dipole in the range $[0,y_{0}]$ can generate with a certain non-zero probability a hadron in the range $[y_{0},y]$. Numerically, the deviation of $C_{0}$ from unity is expected to be about the order of magnitude of the strong coupling constant $\alpha_{s}\simeq 0.2-0.3$. While the original dipole models with fixed dipole sizes provides the KL-model expression Eq. (8) as the correct expression for the inclusive entropy in Eq. (II.2), we will in the following also consider the case where the inclusive entropy is determined by the universal limit Eq. (6).
@@ III Experimental data
To validate the universal approach of describing the QCD evolution of entanglement entropy, the H1 $ep$ DIS data Andreev et al. (2021) from HERA was used. There are two distinct measurements of entanglement entropy in the H1 experiment, see Fig. 2 for an illustration:
@@ V Conclusion
In conclusion, we have studied the QCD evolution of entanglement entropy of the proton in DIS for the first time. We generalized the formalism by Kharzeev and Levin Kharzeev and Levin (2017) to scenarios where final state hadrons are produced in different rapidity intervals. Specifically, our generalized model describes data from both a fixed rapidity window that covers most of the current fragmentation region and a moving, narrow rapidity window centered around the leading parton in DIS. The significant variation in the rapidity intervals of these measurements is sensitive to the QCD evolution of parton distributions.

### v2

@@ I Introduction
Quantum entanglement, famously described by Albert Einstein as a ”spooky action at a distance,” is fundamental to our understanding of the microscopic world. The nontrivial and non-local correlations among the constituents of a quantum system can provide deep insights into the underlying properties of such systems. In recent years, there has been a growing interest in quantum entanglement within high-energy physics, particularly in understanding the structure of hadrons Kharzeev and Levin (2017), thermalization Alba and Calabrese (2017) (see Berges et al. (2021) for a review), spin and helicity coupling of partons Bhattacharya et al. (2024), violation of Bell inequalities in hadronic final states Gong et al. (2022); Aoude et al. (2022); Severi et al. (2022); White and White (2024); CMS (2024); Aad et al. (2023); Banuls et al. (2023), early universe QCD transition Banerjee et al. (2005), and confinement Klebanov et al. (2008); Grieninger et al. (2023a).
@@
The state produced from the proton after the DIS measurement is therefore a mixed state. The corresponding density matrix is obtained by tracing the density matrix of the pure proton state over the unobserved phases of the Fock states, and is characterized by non-zero von Neumann entropy. This entropy can be viewed as the entropy of entanglement between the occupation numbers and the corresponding phases in the Fock state decomposition of the proton wave function. Once this entropy is produced, it will be observed in the final state through the entropy of the multi-hadron state created in DIS. In general, the entanglement entropy may differ from the entropy of the final hadronic state. However there are several examples where the produced entanglement entropy appears equal to the entropy of the final state. One can thus hypothesize that the entanglement entropy is equal to the entropy of the hadron state, and moreover that the hadron multiplicity distribution mirrors the parton occupation number distribution inside the proton. This hypothesis can be viewed as a stronger form of the “local parton-hadron duality”. In previous studies, it has been found that this bold assumption holds surprisingly well Tu et al. (2020); Hentschinski and Kutak (2022); Hentschinski et al. (2022, 2023). Interestingly the model predicts that the entropy linearly grows with rapidity what is in accord with consideration based of near conformal invariance of QCD in the high energy limit Lipatov (1986, 1990, 1994); Braun et al. (2003) and the symmetries of the considered dipole model Gürsoy et al. (2023); Caputa and Kutak (2024).Ref. Dumitru et al. (2023) presents an interesting calculation based on the lowest Fock states of the proton, the 3-quark state and the 3 quark+gluon one. On the other hand, our calculation, as well as the calculation of Kharzeev and Levin (2017), assumes that the maximal entanglement takes place, as a result of a large number of partonic micro-states generated at small x due to QCD evolution. For other works on entanglement entropy at the high energy regime, see Kutak (2011); Peschanski (2013); Armesto et al. (2019); Neill and Waalewijn (2019); Armesto et al. (2019); Kovner et al. (2019); Chachamis et al. (2024); Liu et al. (2022a, 2023a, b, 2023b); Stoffers and Zahed (2013); Asadi and Vaidya (2023); Kou et al. (2022); Kutak (2023); Dumitru et al. (2023); Ehlers (2022, 2023); Florio et al. (2024); Grieninger et al. (2024, 2023b); Ikeda et al. (2023); Brandenburg et al. (2024); Berges et al. (2018a, b); Dumitru and Kolbusz (2023, 2022); Ramos and Machado (2022); Moriggi et al. (2024); Ramos and Machado (2020); Dosch et al. (2024).
@@ II Entanglement entropy in inclusive DIS
In the following we explore both versions of entanglement entropy. Since
@@
is not characteristic for the probability distribution Eq. (2), but also arises as the large $\bar{n}$ limit of other distributions, we refer in the following to it as the “universal entanglement entropy”.
@@
Here $g(x,\mu)$ and $q_{f}(x,\mu)$ denote the PDF of the gluon and the quark of flavor $f$, respectively, which depend on the factorization scale which for the DIS reaction is usually identified with the photon virtuality $Q$. The $\mu$ dependence of PDFs is obtained from the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution equations Gribov and Lipatov (1972); Altarelli and Parisi (1977); Dokshitzer (1977). We note here that in the original K-L model Kharzeev and Levin (2017), only the gluon density was used. As demonstrated in Hentschinski and Kutak (2022), it is indeed the dominant contribution, but for completeness and phenomenological relevance, we also account for quarks.
@@ II.1 Definition of a local dipole probability distribution
The probability to have $n$ dipoles in the interval $[0,y]$ is given by all possible combinations to have $n-m$ dipoles in the range $[0,y_{0}]$ and $m$ in the range $[y_{0},y]$. With
@@ II.2 Definition of local entropy
with a $C_{0}\leq 1$ a constant factor which accounts for subleading effects that modify the no emission probability. With this parameter we will have in particular $\tilde{p}_{0}(y,y)=C_{0}\neq 1$, i.e. the probability for emission, $1-\tilde{p}_{0}$, is non-zero if the rapidity range $y-y_{0}$ shrinks to zero. It therefore introduces a certain non-locality in the parton-hadron duality, which relates the generated dipoles to final state hadrons. While the overall number of dipoles is fixed by the PDFs, a dipole in the range $[0,y_{0}]$ can generate with a certain non-zero probability a hadron in the range $[y_{0},y]$. Numerically, the deviation of $C_{0}$ from unity is expected to be about the order of magnitude of the strong coupling constant $\alpha_{s}\simeq 0.2-0.3$, which characterizes the emission probability of an additional gluon. While the original dipole models with fixed dipole sizes provides the KL-model expression Eq. (8) as the correct expression for the inclusive entropy in Eq. (II.2), we will in the following also consider the case where the inclusive entropy is determined by the universal limit Eq. (6).
@@ III Experimental data
The validation of our model in DIS began by comparing it to inclusive DIS data from H1 Andreev et al. (2021), without considering the finite size of the rapidity window Hentschinski and Kutak (2022); Kharzeev and Levin (2021). We found that the contribution from sea quarks at finite energy was important although not dominant, and incorporating both gluon and quark contributions improved the agreement between the data and the model. Building on this development, we extended the model to describe another QCD process, diffractive DIS, and observed good agreement with our revised model of entanglement entropy Hentschinski et al. (2023). This serves as an important confirmation of our model’s ability to describe two different QCD processes within a consistent framework. However, as previously mentioned, an important aspect that remains to be addressed is the QCD evolution with consideration of varying rapidity windows.
Therefore, to validate the universal approach of describing the QCD evolution of entanglement entropy, including the rapidity window size dependence, the H1 $ep$ DIS data Andreev et al. (2021) from HERA was used. There are two distinct measurements of entanglement entropy in the H1 experiment, see Fig. 2 for an illustration:
@@
Future studies of other QCD processes are also of great interest. For instance, will maximal entanglement be achieved in hard QCD scattering processes, such as events involving high transverse momentum jets and their fragmentation? Could heavy flavor productions offer additional insights into the entanglement-based approach for understanding the parton structure? These questions could be highly relevant for future experimental investigations.
@@ V Conclusion
In conclusion, we have studied the QCD evolution of entanglement entropy of the proton in DIS for the first time. We generalized the formalism by Kharzeev and Levin Kharzeev and Levin (2017) to scenarios where final state hadrons are produced in different rapidity intervals. Specifically, our generalized model describes data from both a fixed rapidity window that covers most of the current fragmentation region and a moving, narrow rapidity window centered around the leading parton in DIS. The significant variation in the rapidity intervals of these measurements is sensitive to the QCD evolution of parton distributions. The new formalism does not require the specific for original K-L model form of the probabilities $P_{n}$. By using different form of $P_{n}$ it can be used to address saturation effects and dipole evolution in a 3+1 D scenario Kharzeev and Levin (2017).